OpenStudy (gentorr):

Solve for x: (-5/4) x - 5 = -35 23 24 −23 −24

OpenStudy (zenmo):

You want to isolate the left side to the number with the "x," so add 5 to both sides.

OpenStudy (gentorr):

so i add -35 with 5

OpenStudy (zenmo):

Yes, what do you get after adding the 5 to both sides?

OpenStudy (en):

its positive 24:)

OpenStudy (gentorr):

ohh olay thanks

OpenStudy (gentorr):

okay

OpenStudy (zenmo):

@En you can't just say the answer, gotta walk her through the process.

OpenStudy (en):

\[(-5/4)x-5=-35\]\[(-5/4)x=-35+5\]

OpenStudy (gentorr):

but thank you any way

OpenStudy (en):

\[(-5/4)x=-30\]\[-5x=-30(4)\]\[-5x=-120\]\[x=24\]

OpenStudy (en):

@GenTorr i posted the formula :) i hope you got the gist of it.

OpenStudy (gentorr):

yes thank you